OpenStudy (anonymous):

A convex decagon has six side equal angles .Two angles are 20 degree more than one of six angles.Two more triangle are 40 degrees more than one of the six angles.Find the measure of each angle.

OpenStudy (anonymous):

total of angles in a decagon = 1440 6x + 2(x+20) + 2(x+40) = 1440 6x + 2x + 2x + 40 + 80 = 1440 10x = = 1320 x =132 2x angles = 264 4xangles = 528

OpenStudy (anonymous):

i assumed you were talking about interior angles

OpenStudy (anonymous):

I'm an idiot Six equal angles of 132 Two equal angles of 152 Two equal angles of 172 ?

OpenStudy (anonymous):

oh yes i worked out other angles incorrectly Elixa has correct solution

OpenStudy (anonymous):

But jimmy's calculations are right if you want working out sorry thanks

OpenStudy (anonymous):

thank Eli and Jimm

OpenStudy (anonymous):

why i doubled and X by 4 I dont know - lol!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Two angles are 20 degree more than one of six angles or each 20 more?

OpenStudy (anonymous):

I think the sum of 2 angle is 20 more not each angle 20 more?

OpenStudy (anonymous):

I take it to mean each of two angles is 20 more than one of the six equall angles

OpenStudy (anonymous):

I was confused when i was working out the value of the two equal angles - i wrote down 2x instead of x + 20

OpenStudy (anonymous):

6x + (x+20) + (x+40) = 1440 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, thank

OpenStudy (anonymous):

six side equal angles, 2 angles (x+20) ,and 2 angles (x+40) ?

OpenStudy (anonymous):

yes